### v1

# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Models are Task-specific Classifiers

Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-source models, especially GPT4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. In this study, we conduct an empirical study of different judge models on their evaluation capability. Our findings indicate that although the fine-tuned judge models achieve high accuracy on in-domain test sets, even surpassing GPT4, they are inherently task-specific classifiers, and their generalizability and fairness severely underperform GPT4.
1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn considerate attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them from comprehensive perspectives Qin et al. (2023). However, traditional evaluation metrics for generative models, such as BLEU and ROUGE, only capture limited aspects of a model’s performance Mathur et al. (2020).
Figure 1: The general training and inference procedure of fine-tuned judge models.
Some research has proposed LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namley utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, LLMs can leverage their instruction-following ability to provide reliable evaluation. For example, Li et al. (2023b) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT4. Their results reveal that strong LLM-based evaluators can achieve a high agreement rate among human experts, establishing a foundation for LLM-as-a-Judge.
Model Foundation Instruction Response Annotation Evaluation Scheme Testset
JudgeLM
(Zhu et al., 2023b) Vicuna Instruct Datasets
(Alpaca-GPT4,
Dolly-15K…) 11 models
(Alpaca,Vicuna…) GPT4 Pairwise Grading GPT4
PandaLM
(Wang et al., 2024) LLaMA Alpaca 52K 5 models
(LLaMA, Bloom…) GPT3.5 Pairwise Selection Human
Auto-J
(Li et al., 2023a) LLaMA2-chat Preference Datasets
(Chatbot Arena,
OpenAI WebGPT…) Preference Datasets Human Pairwise Selection
Pointwise Grading Human
Prometheus
(Kim et al., 2023) LLaMA2-chat GPT4 Generated GPT4 Generated GPT4 Pointwise Grading GPT4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our cross-validation. All the four models are open-source, with their training and test data also publicly released.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed, relying on open-source foundation models and data constructed from either GPT4 or human annotation, as shown in Figure 1. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from GPT4 annotations and fine-tunes a scalable judge model. Auto-J Li et al. (2023a) constructs judgement data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria and finetunes a fine-grained judge model.
In this paper, we conduct an empirical study of the evaluation capabilities of judge models. We conduct extrapolated evaluations among available judge models and meta-evaluation testsets. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations in the following aspects:
  • The fine-tuned judge model is inherently a classification model;
  • The fine-tuned judge model is overfitted to specific evaluation schemes;
  • The fine-tuned judge model is biased towards superficial quality;
To draw a conclusion, the fine-tuned judge models should be used only in similar evaluation scenarios, and can not serve as a general substitution for GPT4 in terms of LLM evaluation.
2 How Far can Fine-tuned Judges Go?
The typical process for finetuning a judge model consists of the following three steps:
Step 1: Data Collection. The training data generally comprises three components: instructions, responses and evaluations. The instructions are typically obtained from instruction datasets, with the responses generated by various representative models, and the evaluations can be derived from either GPT4 or human annotation;
Step 2: Prompt Designing. The prompt template can be structured in various ways depending on the evaluation scheme, such as pairwise selection (which aims to select the better one from a pair of responses), pointwise grading (which aims to score a single reference), etc.
Step 3: Model Fine-tuning. Using the designed prompt and collected data, the training process of the judge model typically follows the instruction fine-tuning paradigm Ouyang et al. (2022). The model is fed with a instruction alongside answer(s) for yielding output comprising evaluation results.
While current judge models mostly self-claim exceeding the evaluation capability of GPT4, in this work, we make an cross validation based on four representative works as listed in Table 1. Our findings are presented in the following sections.
Model Train JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
GPT 3.5 73.83 52.85 62.96 58.20 42.7 0.636 0.563
GPT 4-0613 85.28 76.87 78.68 73.24 56.3 0.742 0.743
[…]
2.1 Inherently a Classification Model
If we do not consider critic generation (which is seldom used in system-level evaluation), the LLM evaluation is inherently a classification (or regression) task. In contrast to the current judge models trained in a generation-style, we train four classification (regression) models based the four groups of data in Table 1111The training settings and prompt templates are presented in Appendix A.1 due to space limitation.. We also train four classification models based on DeBERTaV3-large He et al. (2023) on the same data, which is 20 times smaller than the 7B version of LLaMA.
As shown in Table 2, the classification model performs equally well as the generation model for both pairwise selection and pointwise grading. The powerful generation ability of LLMs hardly brings any improvement to the evaluation accuracy, as they are trained on the same data for the same objective. Moreover, the DeBERTa-based evaluator achieves comparable performance with the LLM-based evaluators [^2] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512 (the maximum context length of DeBERTa)., which might be argued for that the encoder-only architecture is more suitable for classification.
We also analyze the correlation between different predictions made by different evaluators. As shown in Figure 2 and 3, the correlation among different classification models is much closer than their correlation with GPT4. Different as they are in architectures, all three models are inherently classifiers fitting to the same set of supervision, leading to similar evaluation outcomes.
Figure 2: The F1 score between the predictions of different evaluators on JudgeLM testset.
Figure 3: The pearson coefficient between the predictions of different evaluators on Prometheus testset.
2.2 Overfitting to Evaluation Scheme
Model JudgeLM-test PandaLM-test Auto-J-test Average
accuracy F1 accuracy F1 agreement
GPT 3.5 73.83 52.85 62.96 58.20 42.7 59.83
GPT 4-0613 85.28 76.87 78.68 73.24 56.3 73.42
JudgeLM-7B 79.02 71.87 70.97 67.59 46.6 65.53
PandaLM-7B 65.24 47.42 67.57 57.49 40.0 57.61
Auto-J-13B 72.86 57.60 71.47 61.01 54.6 66.31
Prometheus-13B 54.24 50.04 45.25 43.58 47.8 49.10
w/o trans 24.58 23.39 29.03 27.92 16.2 23.26
Table 3: Results of evaluators on pairwise selection. Notice Prometheus can be transformed for pairwise selection by grading two answers twice and compare the scores, therefore we release both results with and without transformation.
Model Prometheus-test-ind Prometheus-test-ood Average
pearson kendalltau spearman pearson kendalltau spearman
GPT 3.5 0.636 0.536 0.617 0.563 0.453 0.521 0.600
GPT 4-0613 0.742 0.659 0.747 0.743 0.660 0.747 0.743
Prometheus-13B 0.864 0.788 0.863 0.869 0.789 0.869 0.867
JudgeLM-7B 0.649 0.647 0.739 0.610 0.602 0.690 0.630
w/o trans 0.398 0.371 0.416 0.384 0.371 0.419 0.391
PandaLM-7B 0.417 0.368 0.423 0.386 0.333 0.383 0.402
Auto-J-13B 0.614 0.526 0.608 0.591 0.504 0.580 0.603
Table 4: Results of evaluators on pointwise grading. Notice JudgeLM can be transformed for pointwise grading by adding the reference as the first answer, therefore we release both results with and without transformation.
Model MTBench
accuracy precision recall F1
GPT 4-0613 66.9 63.8 62.2 61.9
JudgeLM-7B 48.7 52.0 49.7 48.7
PandaLM-7B 55.2 52.6 49.4 46.8
Auto-J-13B 51.7 50.2 46.8 43.7
Prometheus-13B 53.2 49.6 48.4 47.1
Table 5: Results of evaluators on multi-turn evaluation.
One of the most appealing attribute of LLMs is their generalization ability, enabling them to execute various task defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we could verify their evaluation capability under the scheme defined by others. Specifically, we cross-validate the judge models on each other’s testset [^3] We applying their publicly released checkpoints with predefined prompts. For details please refer to Appendix A.2..
As shown in Table 3 and 4, all four models perform the best on their own training settings, respectivly, with results comparable with GPT4. However, if we employ a model on evaluation schemes where it is not trained on [^4] For example, using a pairwise model (such as PandaLM or JudgeLM) for pointwise grading (such as Promethues testset), or using a pointwise model (such as Promethues) for pairwise selection (such as PandaLM or JudgeLM testsets)., the evaluation performance would drop by a large margin. On the contrary, GPT4 consistently exhibits superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset. As can be seen in Table 5, while the four models are all trained for single-turn evaluation, they underperforms GPT4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lack generalibility.
2.3 Biased Towards Superficial Quality
Recently, there has been a few researches on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023); Saito et al. (2023). To address this issue, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and a smooth, coherent but inconsistent answer. We evaluate the judge models on LLMBar\@footnotemark, and the results are shown in Table 6.
It seems that the fine-tuned judge models are severely biased to superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting an accuracy even worse than random guess on the adversarial testsets. On the other hand, GPT4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This also verifies that the fine-tuned judge models are inherently classifiers overfitted to the training data, relying on spurious statistical features for prediction.
It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluator may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation, even if it is not aligned with the instruction.
Model LLMBar
Natu. Neig. GPTI. GPTO. Manu.
GPT 4-0613 93.5 64.2 76.6 76.6 75.0
JudgeLM-7B 62.0 23.1 26.1 46.8 28.3
PandaLM-7B 59.0 16.5 21.7 42.6 26.1
Auto-J-13B 70.0 20.9 21.7 46.8 23.9
Prometheus-7B 53.0 22.4 17.4 27.7 32.6
DeBERTa 62.0 26.9 42.4 55.3 34.8
Table 6: Accuracy of evaluators on bias evaluation.
3 Conclusion
In this work, we empirically study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models underperforms GPT4 in terms of generalibility and fairness by a large margin, which we believe cannot be amended by simply finetuning on more evaluation data. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, watching for the overlap between the evaluation scenario and the training process.
Limitations
Our work still has some limitations: 1) Due to the lack of related work, we primarily relied on cross-validation to assess the generalizability of the four fine-tuned judge models. To conduct a more thorough evaluation of the generalizability, it would be beneficial to incorporate additional independent testsets encompassing a broader range of evaluation schemes, such as reference augmentation and domain-specific evaluation. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as formality bias, verbosity bias, etc. 3) Due to time and resource constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
Appendix A Appendix
A.1 Training Settings
As mentioned in Section 2.1, we fine-tune our own judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison. We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyper-parameters are presented in Table 7. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lies in the prediction method. For generation model, the prediction head reused the pretrained language model head, and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^5] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details., as shown in Figure 4.
A.2 Prompt Templates
As mentioned in Section 2.1, 2.2 and 2.3, we take the publicly released checkpoints of the four models and validate their performance. We use the same prompt templates as defined by the four open-source models, as presented from Figure 5 to Figure 12. For JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading.
Configuration Vicuna-based DeBERTa-based
base model Vicuna-7B DeBERTaV3-large
max length 2048 512
learning rate 2e-5 2e-5
learning rate schedule cosine decay cosine decay
optimizer AdamW AdamW
AdamW beta1 0.9 0.9
AdamW beta2 0.999 0.98
weight decay 0.0 0.0
GPU nums 8 2
training epochs 3 3
batch size 128 128
warmup ratio 0.003 0.003
numerical precision bf16, tf32 fp16
ZeRO optimizer stage 2 None
Table 7: Configurations of the judge models fine-tuned by us in Section 2.1. Both classification and generation judge models leverage the same group of configs based on their foundation model.
Figure 4: The architecture of classification-based judge model. The major difference lies in the prediction head, where a new classification (regression) head is initialized for predicting the result.
Figure 5: Prompt template for JudgeLM applied for pairwise selection.
Figure 6: Prompt template for JudgeLM applied for pointwise grading.
Figure 7: Prompt template for PandaLM applied for pairwise selection.
Figure 8: Prompt template for PandaLM applied for pointwise grading.
Figure 9: Prompt template for Auto-J applied for pairwise selection.
Figure 10: Prompt template for Auto-J applied for pointwise grading.
Figure 11: Prompt template for Prometheus applied for pairwise selection.
Figure 12: Prompt template for Prometheus applied for pointwise grading.