### v4

# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Model is not a General Substitute for GPT-4

Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have fine-tuned judge models based on open-source LLMs for evaluation. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this work, we conduct an empirical study of LLM-as-a-Judge. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness and adaptability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations1.
footnotetext: * Equal contribution. ✉ Corresponding Author.footnotetext: 1Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge.
1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn significant attention Liang et al. (2022); Chang et al. (2023); He et al. (2024); Gu et al. (2025); He et al. (2025). Some research has proposed LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, proprietary LLMs can provide an accurate evaluation with high agreement with human evaluators.
Figure 1: The general training and inference procedure of fine-tuned judge models.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2024); Wang et al. (2024); Ke et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
{tblr}
hlines, Model & Foundation Instruction Response Annotation Evaluation Scheme Testset
JudgeLM
(Zhu et al., 2024) Vicuna Instruct Datasets
(Alpaca-GPT4,
Dolly-15K…) 11 models
(Alpaca,Vicuna…) GPT-4 Pairwise Grading GPT-4
PandaLM
(Wang et al., 2024) LLaMA Alpaca 52K 5 models
(LLaMA, Bloom…) GPT3.5 Pairwise Selection Human
Auto-J
(Li et al., 2024a) LLaMA2-chat Preference Datasets
(Chatbot Arena,
OpenAI WebGPT…) Preference Datasets Human Pairwise Selection
Pointwise Grading Human
Prometheus
(Kim et al., 2024) LLaMA2-chat GPT-4 Generated GPT-4 Generated GPT-4 Pointwise Grading GPT-4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test MT-Bench
accuracy F1 accuracy F1 agreement PCC-ind PCC-ood accuracy F1
JudgeLM-7B 82.39 72.97 68.17 65.18 45.3 0.398 0.384 48.7 48.7
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 0.417 0.386 55.2 46.8
Auto-J-13B 77.79 62.64 72.17 64.10 53.6 0.614 0.591 51.7 43.7
Prometheus-13B 24.58 23.39 29.03 27.92 16.2 0.864 0.869 53.2 47.1
+grade-twice 54.24 50.04 45.25 43.58 47.8
Deepseek-V3 79.23 68.27 75.97 71.25 57.0 0.734 0.741
GPT-4-mini 79.17 68.31 76.57 71.79 57.4 0.707 0.705
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106 84.24 72.83 75.78 71.51 56.9 0.742 0.743 66.9 61.9
Table 2: Results of evaluators on different evaluation schemes. Notice JudgeLM-test, PandaLM-test, Auto-J-test are pairwise selection, Prometheus-test is pointwise grading, and MT-Bench is multi-turn evaluation.
In this paper, we conduct an empirical study for the evaluation capability of judge models. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced proprietary models:
  • The fine-tuned judge model is constrained by specific evaluation scheme;
  • The fine-tuned judge model is biased towards superficial quality;
  • The fine-tuned judge model is incapable of aspect-specific evaluation;
  • The fine-tuned judge model can not benefit from prompting strategies;
We argue that these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data. To draw a conclusion, the fine-tuned judge model cannot serve as a general substitute for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging them for evaluation in real applications, watching for the overlap between the evaluation scenario and the fine-tuning process.
2 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models in Table 1111We make minimal change to the predefined prompts to adapt the judge model to different schemes. Please refer to Appendix A.2 for detailed implementations., and reveal there exist several limitations about their evaluation capabilities.
2.1 Constrained by Evaluation Scheme
One of the most appealing attributes of LLMs is their generalization ability, enabling them to execute various tasks defined by various instructions Zhu et al. (2023). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, chain-of-thought evaluation, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their capability on uncovered schemes. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets. We also validate the models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset.
HaluEval-QA HaluEval-Sum HaluEval-Dial ToxicChat SALAD-Bench
Model accuracy F1 accuracy F1 accuracy F1 accuracy F1 accuracy F1
JudgeLM-7B - - - - - - - - 82.45 57.44
PandaLM-B - - - - - - - - 57.03 37.23
Auto-J-13B 58.30 56.03 53.10 43.34 63.10 62.90 87.40 52.24 86.88 52.66
w/o adapt 59.60 57.38 53.47 43.55 64.50 63.71 87.70 51.15 71.77 47.86
Prometheus-7B 47.90 45.84 44.50 40.38 51.00 45.17 77.10 58.14 - -
w/o adapt 48.90 45.10 46.60 36.43 53.40 50.24 81.20 61.87 - -
GPT-3.5-0613 57.50 57.10 62.60 60.27 72.10 72.08 95.10 80.80 95.54 61.70
GPT-4-1106 72.50 72.50 72.00 71.44 84.50 84.78 94.50 82.78 98.75 65.55
Table 3: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. For more details please refer to A.2.
Model LLMBar
Natu. Neig. GPTI. GPTO. Manu.
JudgeLM-7B 62.0 23.1 26.1 46.8 28.3
PandaLM-7B 59.0 16.5 21.7 42.6 26.1
Auto-J-13B 70.0 20.9 21.7 46.8 23.9
Prometheus-7B 53.0 22.4 17.4 27.7 32.6
GPT-4-1106 93.5 64.2 76.6 76.6 75.0
Table 4: Accuracy of evaluators on bias evaluation.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement PCC-ind PCC-ood
Released Models 82.39 72.97 68.97 60.95 53.6 0.864 0.869
Vicuna-generation 82.44 71.77 72.37 60.78 47.6 0.826 0.815
Vicuna-classification 82.16 70.07 70.87 60.34 46.8 0.846 0.831
DeBERTa-classification 81.30 68.34 72.27 51.75 31.7 0.835 0.813
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106-preview 84.24 72.83 75.78 71.51 56.9 0.742 0.743
Table 5: Comparison of generation and classification-based evaluators. Results with are from evaluating the four publicly released models on their respective testsets, and results with are from evaluating models trained by us.
As shown in Table 2, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained, the evaluation performance would drop by a large margin. On the contrary, close-sourced proprietary models such as GPT-3.5 or GPT-4 consistently exhibit superior performance across various evaluation schemes.
Model Method JudgeLM-test PandaLM-test Auto-J-test Salad-bench
accuracy F1 acuracy F1 agreement accuracy F1
JudgeLM-7B w/o CoT 82.74 72.64 72.67 69.32 44.25 85.57 57.35
w/ original CoT 82.26 67.41 70.77 64.24 41.45 75.15 50.59
w/ o1 CoT 77.96 65.43 66.46 62.95 37.90 72.39 47.91
Table 6: Comparison of different CoT sources on JudgeLM-7B.
JudgeLM-test PandaLM-test
Model accuracy F1 accuracy F1
JudgeLM-7B 82.39 72.97 68.17 65.18
+ CoT 81.68 71.59 68.03 64.42
+ ICL 68.57 58.52 41.14 40.39
PandaLM-7B 66.44 56.01 68.97 60.95
+ CoT 65.85 56.59 68.03 60.42
+ ICL 66.16 55.94 68.97 59.40
Auto-J-13B 77.79 62.64 72.17 64.10
+ ICL 76.20 59.12 68.37 58.44
GPT-3.5-0613 72.57 51.40 64.36 46.40
+ CoT 75.24 60.71 69.97 63.66
+ ICL 69.38 57.46 70.67 56.12
GPT-4-1106 84.24 72.83 75.78 71.51
+ CoT - - 77.08 71.77
+ ICL - - 64.86 56.20
Table 7: Results of evaluators with ICL and CoT. We did not apply GPT-4 on JudgeLM-test as the annotation of JudgeLM-test is conducted with GPT-4 without ICL and CoT. We only apply ICL on Auto-J as the original prompt of Auto-J comprises CoT.
2.2 Biased Towards Superficial Quality
Recently, there has been a lot of research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). Subsequently, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises four adversarial testsets (Neig., Manu., GPTO., GPTI.) with paired outputs of a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.).
We evaluate the judge models on LLMBar. As shown in Table 4, the fine-tuned judge models perform poorly on adversarial testsets, even worse than random-guess. This notifies that they are severely biased toward superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference for incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on LLMBar.
2.3 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as helpfulness, safety, etc. In this part, we would like to assess the evaluation capability of judge models on fine-grained aspects, based on the following datasets: 1) HaluEval Li et al. (2023a): for factuality evaluation; 2) ToxicChat Lin et al. (2023): for toxicity evaluation; 3) SALAD-Bench Li et al. (2024b): for safety evaluation.
As can be seen from Table 3, the fine-tuned judges fall far behind on all fine-grained aspects. It deserves to notice that while Prometheus is designed for fine-grained evaluation, it obtains an inferior performance on both benchmarks, which notifies that it failed to learn the correlation between fine-grained aspects and evaluation results.
For the purpose of comparison, we also apply Auto-J and Prometheus with their original prompt on aspect-specific evaluation. As can be seen in Table 3, to our surprise, their performance remains roughly the same compared with aspect-specific prompts, notifying that both models have lost the general instruction-understanding ability, therefore the aspect-specific prompt is not taking effect.
2.4 Can not Benefit from CoT and ICL
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies, namely In-context Learning (ICL) Dong et al. (2023) and Chain-of-Thought Prompting (CoT) Wei et al. (2022), to further improve the evaluation capability of the judge models.
As shown in Table 7, while the close-sourced proprietary models are improved by a large margin through both prompt engineering strategies, the fine-tuned judges hardly benefit from these strategies, sometimes even experiencing severe performance decline. Specifically, in the case of CoT prompting, despite we modified the prompts for JudgeLM and PandaLM to generate CoT firstly, both models failed to produce CoT and adhered to their original output format, as they have lost their general instruction-following ability.
We also evaluated the impact of different CoT sources based on JudgeLM-7B. We first swapped the positions of scores and CoT in the training data, and then fine-tuned the base model with or without CoT using the same hyperparameters. Additionally, we utilized o1-preview-0912222platform.openai.com/docs/models/o1 to generate CoT for the original scores through hint-driven prompting Srivastava et al. (2023), and subsequently fine-tuned the model with this annotated CoT.
As demonstrated in Table 6, fine-tuning with either the original CoT or the o1-generated CoT resulted in a degradation of model performance compared to the model fine-tuned without CoT. Notably, the o1-generated CoT led to a more severe performance drop. This clearly indicates that even high-quality CoT did not introduce any improvement to the fine-tuned judge.
3 The Essence of Fine-tuned Judge: A Task-specific Classifier
Figure 2: The F1 score between the predictions of different evaluators on JudgeLM testset.
Figure 3: The pearson coefficient between the predictions of different evaluators on Prometheus testset.
Combining all the limitations revealed in our experiments, we would like to claim that after the fine-tuning process on a single task, the judge model has degenerated into a task-specific classifier, which is overfitted to the training data. To support this, we fine-tune three groups of judges based on the four groups of data as listed in Table 1333Please refer to Appendix A.1 for training details.:
  1. 1.
    Vicuna-generation Chiang et al. (2023): It formulates the evaluation task in a generation-style, and the prediction head reuses the pretrained language model head;
  2. 2.
    Vicuna-classification: It formulates the evaluation task as classification or regression, and the prediction head is newly initialized as a linear projection layer;
  3. 3.
    DeBERTa-classification: It also formulates as a classification task, based on DeBERTaV3-large He et al. (2023), which is 20 times smaller than the 7B version of Vicuna;
As shown in Table 5, the classification model performs equally well as the generation model. The formidable generative capabilities of LLMs hardly bring any improvement to the evaluation, as they are fitting to the same group of data. Moreover, the DeBERTa-based classifier achieves comparable performance with the LLM-based evaluators [^4] The only exception is on Auto-J-test, which is possibly due to a large proportion of the test data exceeds 512., which might be argued for that the encoder-only architecture is more suitable for classification.
We also analyze the correlation between different predictions made by different evaluators. As shown in Figure 2 and 3, the correlation among different classification models is much closer than their correlation with GPT-4. Different as they are in architectures, all three models are inherently classifiers fitting to the same set of supervision, leading to similar evaluation outcomes.
Although prior research on instruction-tuning all emphasizes the importance of data diversity Zhou et al. (2023); Lu et al. (2024), the fine-tuning of judges is doing the opposite thing. Therefore, after fine-tuning for a single task with a fixed prompt template, the model lost its generalization ability, and degenerate into a task-specific classifier, which exhibits several limitations due to overfitting.
4 Conclusion
Although the fine-tuned models demonstrate superior performance on in-domain test sets, they still have several limitations compared to GPT-4. While increasing the fine-tuning data could possibly mitigate some of the limitations, as the potential of LLM extends beyond boundaries, there will always be new domains and tasks that are not covered by the fine-tuning scope. Therefore, the fine-tuned judge model cannot replace GPT-4 as a universal evaluator for LLMs, and should be used judiciously by watching the domain and task adaptability.
Limitations
Our work still has some limitations: 1) Due to time limitation, we did not present a possible solution to mitigate the limitations of fine-tuned judge models. We will investigate related method in the future. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
Acknowledgements
This work is supported by National Natural Science Foundation of China (62276077, 62376075, 62376076).
Appendix A Appendix
A.1 Training Settings
As mentioned in Section 1, we fine-tune our judge models based on the four groups of data (JudgeLM Zhu et al. (2024), PandaLM Wang et al. (2024), Auto-J Li et al. (2024a), Prometheus Kim et al. (2024)), both in generation-style and in classification-style, for the purpose of comparison.
Configuration Vicuna DeBERTa
max length 2048 512
learning rate 2e-5 2e-5
scheduler cosine decay cosine decay
optimizer AdamW AdamW
AdamW beta1 0.9 0.9
AdamW beta2 0.999 0.98
weight decay 0.0 0.0
training epochs 3 3
batch size 128 128
warmup ratio 0.003 0.003
numerical precision bf16 fp16
ZeRO optimizer stage 2 None
Table 8: Configurations of the fine-tuned judge models. Both classification and generation models leverage the same group of configs based on their foundation model.
We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyperparameters are presented in Table 8. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lying in the prediction method, as illustrated in Figure 4. For generation model, the prediction head reused the pretrained language model head and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^5] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details..
Figure 4: The architecture of classification-based judge model. The major difference lies in the prediction head, where a new classification (regression) head is initialized for predicting the result.
A.2 Prompt Templates
As mentioned in Section 1, we take the publicly released checkpoints of the four fine-tuned judge models and validate their performance. To make a fair comparison, we make minimal modifications to their pre-defined prompts, to adapt them to different scenarios, as listed as follows.
For Section 2.1, we adopt the prompts presented in Figure 5 to 12 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 13 to 16 on MT-Bench, which are all adapted to multi-turn evaluation. We adopt the prompts presented in Figure 21 and Figure 22 for chain-of-thought prompting.
For Section 2.2, we adopt the prompts presented in Figure 5, 7, 9 and 11, as LLMBar is in the form of pair-wise selection.
For Section 2.3, we adopt the prompts presented in Figure 17 to 20 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks. As HaluEval and ToxicChat are both binary classifications, we apply Auto-J and Prometheus with pointwise grading and conduct a grid search to determine the classification threshold. On the other hand, as SALAD-Bench is a pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM, and Auto-J to select a better response.
Figure 5: Prompt template for JudgeLM applied for pairwise selection.
Figure 6: Prompt template for JudgeLM applied for pointwise grading.
Figure 7: Prompt template for PandaLM applied for pairwise selection.
Figure 8: Prompt template for PandaLM applied for pointwise grading.
Figure 9: Prompt template for Auto-J applied for pairwise selection.
Figure 10: Prompt template for Auto-J applied for pointwise grading.
Figure 11: Prompt template for Prometheus applied for pairwise selection.
Figure 12: Prompt template for Prometheus applied for pointwise grading.
Figure 13: Prompt template for JudgeLM applied for multi-turn grading.
Figure 14: Prompt template for PandaLM applied for multi-turn grading.
Figure 15: Prompt template for Auto-J applied for multi-turn grading.
Figure 16: Prompt template for Prometheus applied for multi-turn grading.
Figure 17: Prompt template for JudgeLM applied on SALAD-Bench.
Figure 18: Prompt template for Auto-J applied on HaluEval.
Figure 19: Prompt template for Auto-J applied on ToxicChat.
Figure 20: Prompt template for Auto-J applied on SALAD-Bench.
Figure 21: Prompt template for JudgeLM applied with chain-of-thought prompting.
Figure 22: Prompt template for PandaLM applied with chain-of-thought prompting.